# Inner-Probe: Discovering Copyright-related Data Generation in LLM Architecture

### v2

@@ IV-D Controllable Text Generation Design
Copyright law uses the ‘substantial similarity’ criteria to identify infringement in books. This criterion examines similarities in unique expressions, plot elements, character relationships, and settings between works [50].
Based on this, we design a controlled generation process for books using LLMs’ language understanding ability. By incorporating these criteria into prompt templates and providing copyrighted source text fragments, we guide LLMs to generate texts with deterministic copyright-infringement to the original works. To extract relevant text fragments from copyrighted authors, we use the Kimi API from Moonshot AI [51], with a complete process shown in Fig. 8.
@@
Criteria for books cannot be directly applied to other datasets. Therefore, to extend to broader datasets in the Pile [23], we design three controlled text generation scenes: continuation, copyright-transfer, and copyright-elimination (Fig. 9). Each scene represents a distinct way of generating potentially copyright-related content.
@@
V Case Study: Attributing Real-world Copyright-infringed LLM Generated Texts
The Books3 subset in the Pile [23] has been proven to contain copyrighted books from famous authors and used for LLM training, such as GPT-Neo and GPT-J [52, 53]. Our task is to determine which author contributes the most if the generated text has a copyright violation. In this section, we show Inner-Probe is more effective in such analysis in real-world scenarios than other methods.
TABLE I: Copyrighted books identified in Books3 dataset for copyright analysis in case study.
@@ V-A Experiment Setup
We choose eight books from representative copyrighted authors, [54], which are included in Books3 [23]. The full list is shown in Tab. I. We focus on developing a method that can effectively work with limited training data, which is crucial for real-world copyright protection scenarios where extensive training samples from each author may not be readily available or practical to obtain. Thus, we randomly extract 160 snippets of 512 tokens from author’s works for training, and 4800 snippets for validation. This formulates an eight-class classification problem where the model learns to distinguish between authors based on their original copyrighted texts.
To evaluate Inner-Probe’s ability to classify in real-world copyright-infringed texts, we construct a test set by selecting representative text snippets, and then prompt target LLM for text generation. We first select 400 text snippets that embody their distinctive themes or writing styles using the Kimi API [51]. These snippets are then processed in target LLM using our prompt engineering method (sec. IV-D) to generate infringed texts. The model’s classification accuracy on these generated samples directly measures its capability to trace LLM-generated texts back to the original authors. Specifically, higher accuracy indicates better performance in identifying which author’s work has been infringed by the LLM.
We also conduct non-copyright response filtering task, where 900 non-copyrighted texts are collected from Gutenberg [55], and 900 copyrighted texts from Books3.
@@
We propose two baselines for the copyright sub-datasets contribution analysis. The prompt engineering baseline ($Baseline_{1}$/B1) is designed based on templates in [56], treating LLMs as agents to ask them to decide the sub-dataset contributions according to the given texts. For $Baseline_{2}$ (B2), we implement a distance-based approach to show the importance of layer-wise dependencies. This method computes class-specific representative embeddings by averaging MHA across all layers of training samples within each class. For classification, the test sample is assigned to the class whose representative has the shortest L1 distance to the sample’s embedding. Thus, $Baseline2$ uses LLM architecture but ignores layer-wise patterns.
Besides prompt-based attribution and distance-based methods, text classification methods (BERT [57] and RoBERTa [58] both with full parameter finetuning) are compared simultaneously.
V-B Results and Analysis
We evaluate all methods on both copyrighted books and LLM-generated texts. For supervised methods, models are trained on copyrighted books. Meanwhile, prompt-based attribution directly queries target LLM by prompt templates. On the evaluation of copyrighted books, standard text classification methods and our proposed methods both achieve over 95% accuracy. The remaining methods (prompt attribution and distance-based methods) perform significantly worse, all below 40%, as shown in Tab. II.
To further distinguish text classification methods and Inner-Probe, we notice when evaluating on LLM-generated copyrighted texts, Inner-Probe (shown in bold in the 3nd column in Tab. II) consistently outperforms other methods by over 8.0% in accuracy. This is because Inner-Probe effectively utilizes MHA features, which better correlates the distributional shifts between original copyrighted books and LLM-generate texts. We also find our method performs well on non-copyright response filtering in Tab. III.
TABLE II: Classification accuracy of different methods in the case study. GPT-Neo-2.7B generates copyright-infringed texts, and serves as target LLM for Inner-Probe, prompt attribution, and distance-based methods.
Method Accuracy(%$\uparrow$ Original Book Accuracy(%$\uparrow$ Generated Text Cost(Min/GB)$\downarrow$ Time/Memory
@@ V-B Results and Analysis
BERT [57] 98.1 72.7 29.4/58.3
RoBERTa [58] 96.1 72.5 29.5/58.8
@@
TABLE III: Non-copyright Dataset Detection Accuracy and AUC Comparison on GPT-Neo-2.7B in the case study.
Method Inter. Var. A-Var. Baseline [21]
@@
VI Experiment
VI-A Experiment Setup
@@
We test our methods over three open-source language models, including encoder-based and decoder-based ones: bert-base-uncased [59], GPT-Neo-2.7B [52] and GPT-J-6B [53].
For the encoder-only model, we evaluate six GLUE classification sub-datasets: CoLA, SST-2, MRPC, WNLI, QQP, and RTE, with each sub-dataset comprising 500 training and 100 testing samples. Before further experiments, BERT is fine-tuned on tasks in GLUE. For the decoder-based models, experiments utilize the Pile dataset, an 825 GB open-source collection of 22 diverse sub-datasets [23]. 8 classes including Github, OpenWebText2, Wikipedia (en), StackExchange, PubMed Abstracts, Pile-CC, USPTO Backgrounds, and FreeLaw are considered copyrighted datasets, and 4 classes including PubMed Central, Enron Emails, OpenSubtitles, and DM Mathematics are considered non-copyrighted ones.
@@
For the copyright sub-datasets contribution analysis module, we compare our method with text classification models, including BERT [57], and RoBERTa [58]. We also implement a prompt-based method ($Baseline1$/B1) and a distance-based method ($Baseline2$/B2).
For the non-copyrighted responses filtering module, we compare with the unsupervised method [21] as the baseline. We evaluate it under two settings for fair comparisons: (Setting1) with copyright sub-category information incorporated through an extra cross-entropy loss implemented in [21]. (Setting2) treating all copyright data as a single class in unsupervised learning. Our method uses Setting2, and both settings are used in the baseline.
VI-B Copyright Sub-datasets Contribution Analysis Module
VI-B1 Classification Performance on Original Copyrighted Texts
@@
VI-B2 Evaluation on Controlled Generation Scenes
@@
We conduct experiments on GPT-Neo-2.7B and GPT-J-6B, and find that $K_{IB}$ accurately predicts the classification accuracy. As depicted in Tab. IV, $K_{IB}$ correlates with classification accuracy with exactly the same trend. The interval-based method shows better IB metrics, which corresponds to its higher accuracy.
@@
TABLE IV: Explaining the superior performance of Inter in Inner-Probe using Mutual Information analysis on GPT-Neo-2.7B. Notably, the $K_{IB}$ results demonstrate a consistent trend with classification accuracy. Results of GPT-J-6B are not shown for space limit.
@@
VI-C Non-copyrighted Responses Filtering Module
@@
Incorporating intra-class copyright information improves performance, but using LSTM prior learnt from LLM has more to offer compared with using it during the unsupervised learning stage. As in Tab. V, baseline [21] with Setting1 improves accuracy by 3.94% and AUC by 0.108, yet still underperforms our method (Setting2) by an AUC of 0.104. Our method (Inter, Var and A-Var) shows consistent advantages and is more computationally efficient, as it doesn’t require fine-tuning on language models like BERT.
TABLE V: Non-copyright Detection Accuracy and AUC Results.
Inter. Var. A-Var. Baseline [21]
@@
VI-D Training Efficiency and Data Scale Effect
Robustness to Data Scale Large amounts of copyrighted data pose challenges to model training, necessitating a lightweight method. To evaluate data scaling effects, we test Inner-Probe by setting maximum token number $k$ to 7, and varying training samples from 160 to 1,600 and 12,800, with 6,400 inference samples. As shown in Tab. VI, training on 1,600 samples is sufficient for Inner-Probe to outperform baselines, while reducing training time (Tab. VII). The accuracy increases linearly with data scale and $k$. Our empirical study shows that while a small $k$ is generally sufficient, $k=7$ leads to model degradation.
TABLE VI: Data scale effect across different methods and models. $k$ denotes the sampling token number. The model accuracy strictly follows the increase in $k$ and data scale until $k=7$.
@@ VI-D Training Efficiency and Data Scale Effect
Small (160,6400) 1 69.6 75.6 63.4 75.7 75.4 70.8 - - - -
3 85.5 87.1 76.9 83.5 83.8 80.4 - - - -
5 85.9 88.6 79.7 85.9 85.9 82.7 - - - -
7 82.7 87.6 67.4 84.7 79.0 81.4 - - 39.2 72.3
Medium (1600,6400) 1 88.6 89.0 89.9 91.2 89.5 88.3 - - - -
3 92.8 93.1 91.5 92.6 91.8 91.1 - - - -
5 93.3 93.6 92.3 93.2 93.1 91.3 - - - -
7 93.4 94.0 92.5 93.2 93.0 91.7 - - 42.7 79.4
Large (12800,6400) 1 92.0 93.3 93.4 93.9 92.2 93.3 - - - -
3 94.8 95.4 94.3 95.3 94.3 94.4 - - - -
5 95.4 95.5 94.3 95.4 94.9 94.7 - - - -
7 95.2 95.7 94.8 95.3 95.2 95.1 14.0 15.6 42.8 79.9
Training Efficiency We conduct a comparative analysis of time cost across different methods and models. As shown in Tab. VII, Inner-Probe shows superior performance in less than 1.5 hours, far negligible than training or finetuning LLMs [60].
TABLE VII: Training Efficiency Comparison
@@
VII Conclusion
We highlight the need to identify specific copyrighted sub-datasets’ influence on LLM outputs. Our proposed framework Inner-Probe looks into LLM architectures, and discovers that MHA outputs provide more effective information than prompt-based, distance-based, or text classification methods. Information extraction, copyright sub-datasets contribution analysis and non-copyright responses filter modules are designed. Inner-Probe shows high accuracy and AUC comparable to text classification baselines when tested on original copyrighted texts. We further validate various generation scenarios, including prompt-induced book copyright infringements, continuation, copyright transfer, copyright elimination, and text mixture. Inner-Probe efficiently combines multi-modules for copyrighted data protection with minimal data, time, and GPU memory need, outperforming text classification, prompt-based and distance-based baselines by more than 15.02% accuracy and 0.104 AUC increase.

### v3

@@ 4.4 Controllable Text Generation Design
Copyright law uses the ‘substantial similarity’ criteria to identify infringement in books. This criterion examines similarities in unique expressions, plot elements, character relationships, and settings between works [wikipediaSubstantialSimilarity].
Based on this, we design a controlled generation process for books using LLMs’ language understanding ability. By incorporating these criteria into prompt templates and providing copyrighted source text fragments, we guide LLMs to generate texts with deterministic copyright-infringement to the original works. To extract relevant text fragments from copyrighted authors, we use the Kimi API from Moonshot AI [Qin2024MooncakeAK], with a complete process shown in Fig. 8.
@@
Criteria for books cannot be directly applied to other datasets. Therefore, to extend to broader datasets in the Pile [gao2020pile], we design three controlled text generation scenes: continuation, copyright-transfer, and copyright-elimination (Fig. 9). Each scene represents a distinct way of generating potentially copyright-related content.
@@
5 Case Study: Attributing Real-world Copyright-infringed LLM Generated Texts
The Books3 subset in the Pile [gao2020pile] has been proven to contain copyrighted books from famous authors and used for LLM training, such as GPT-Neo and GPT-J [Neo, gpt-j]. Our task is to determine which author contributes the most if the generated text has a copyright violation. In this section, we show Inner-Probe is more effective in such analysis in real-world scenarios than other methods.
Table 1: Copyrighted books identified in Books3 dataset for copyright analysis in case study.
@@ 5.1 Experiment Setup
We choose eight books from representative copyrighted authors, [theguardianZadieSmith], which are included in Books3 [gao2020pile]. The full list is shown in Tab. 1. We focus on developing a method that can effectively work with limited training data, which is crucial for real-world copyright protection scenarios where extensive training samples from each author may not be readily available or practical to obtain. Thus, we randomly extract 160 snippets of 512 tokens from author’s works for training, and 4800 snippets for validation. This formulates an eight-class classification problem where the model learns to distinguish between authors based on their original copyrighted texts.
To evaluate Inner-Probe’s ability to classify in real-world copyright-infringed texts, we construct a test set by selecting representative text snippets, and then prompt target LLM for text generation. We first select 400 text snippets that embody their distinctive themes or writing styles using the Kimi API [Qin2024MooncakeAK]. These snippets are then processed in target LLM using our prompt engineering method (sec. 4.4) to generate infringed texts. The model’s classification accuracy on these generated samples directly measures its capability to trace LLM-generated texts back to the original authors. Specifically, higher accuracy indicates better performance in identifying which author’s work has been infringed by the LLM.
We also conduct non-copyright response filtering task, where 900 non-copyrighted texts are collected from Gutenberg [gutenbergProjectGutenberg], and 900 copyrighted texts from Books3.
@@
We propose two baselines for the copyright sub-datasets contribution analysis. The prompt engineering baseline ($Baseline_{1}$/B1) is designed based on templates in [person], treating LLMs as agents to ask them to decide the sub-dataset contributions according to the given texts. For $Baseline_{2}$ (B2), we implement a distance-based approach to show the importance of layer-wise dependencies. This method computes class-specific representative embeddings by averaging MHA across all layers of training samples within each class. For classification, the test sample is assigned to the class whose representative has the shortest L1 distance to the sample’s embedding. Thus, $Baseline2$ uses LLM architecture but ignores layer-wise patterns.
Besides prompt-based attribution and distance-based methods, text classification methods (BERT [kenton2019bert] and RoBERTa [liu2019roberta] both with full parameter finetuning) are compared simultaneously.
5.2 Results and Analysis
We evaluate all methods on both copyrighted books and LLM-generated texts. For supervised methods, models are trained on copyrighted books. Meanwhile, prompt-based attribution directly queries target LLM by prompt templates. On the evaluation of copyrighted books, standard text classification methods and our proposed methods both achieve over 95% accuracy. The remaining methods (prompt attribution and distance-based methods) perform significantly worse, all below 40%, as shown in Tab. 2.
To further distinguish text classification methods and Inner-Probe, we notice when evaluating on LLM-generated copyrighted texts, Inner-Probe (shown in bold in the 3nd column in Tab. 2) consistently outperforms other methods by over 8.0% in accuracy. This is because Inner-Probe effectively utilizes MHA features, which better correlates the distributional shifts between original copyrighted books and LLM-generate texts. We also find our method performs well on non-copyright response filtering in Tab. 3.
Table 2: Classification accuracy of different methods in the case study. GPT-Neo-2.7B generates copyright-infringed texts, and serves as target LLM for Inner-Probe, prompt attribution, and distance-based methods.
[…]
@@ 5.2 Results and Analysis
BERT [kenton2019bert] 98.1 72.7 29.4/58.3
RoBERTa [liu2019roberta] 96.1 72.5 29.5/58.8
@@
Table 3: Non-copyright Dataset Detection Accuracy and AUC Comparison on GPT-Neo-2.7B in the case study.
Method Inter. Var. A-Var. Baseline [cho2022enhancing]
@@
6 Experiment
6.1 Overview
Following the Books3 dataset in the case study, we first evaluate Inner-Probe on its parent corpus, the Pile [gao2020pile], in the main experiments. These datasets are widely used and known for copyright issues in LlaMA and GPT pretraining [touvron2023llama, detecting, das2025blind, tao2025evidencing], yet they offer limited coverage and only partial transparency for disclosed LLMs. To expand coverage, we add analysis on the larger, multilingual dataset called Matrix. To improve transparency, we use an open-source LLM with a public pretraining pipeline and dataset, namely Map-Neo 7B trained on Matrix [zhang2024map].
Therefore, we report results in two tracks. Besides the main experiments, Section 6.6 first extends experiments to the Matrix dataset, then discusses limitations by studying how LLM generation length and data scale affect Inner-Probe.
6.2 Experiment Setup
@@
We test our methods over three open-source language models, including encoder-based and decoder-based ones: bert-base-uncased [bert-base-uncased], GPT-Neo-2.7B [Neo] and GPT-J-6B [gpt-j].
For the encoder-only model, we evaluate six GLUE classification sub-datasets: CoLA, SST-2, MRPC, WNLI, QQP, and RTE, with each sub-dataset comprising 500 training and 100 testing samples. Before further experiments, BERT is fine-tuned on tasks in GLUE. For the decoder-based models, experiments utilize the Pile dataset, an 825 GB open-source collection of 22 diverse sub-datasets [gao2020pile]. 8 classes including Github, OpenWebText2, Wikipedia (en), StackExchange, PubMed Abstracts, Pile-CC, USPTO Backgrounds, and FreeLaw are considered copyrighted datasets, and 4 classes including PubMed Central, Enron Emails, OpenSubtitles, and DM Mathematics are considered non-copyrighted ones.
@@
For the copyright sub-datasets contribution analysis module, we compare our method with text classification models, including BERT [kenton2019bert], and RoBERTa [liu2019roberta]. We also implement a prompt-based method (Baseline1Baseline1/B1) and a distance-based method (Baseline2Baseline2/B2).
For the non-copyrighted responses filtering module, we compare with the unsupervised method [cho2022enhancing] as the baseline. We evaluate it under two settings for fair comparisons: (Setting1) with copyright sub-category information incorporated through an extra cross-entropy loss implemented in [cho2022enhancing]. (Setting2) treating all copyright data as a single class in unsupervised learning. Our method uses Setting2, and both settings are used in the baseline.
6.3 Copyright Sub-datasets Contribution Analysis Module
6.3.1 Classification Performance on Original Copyrighted Texts
@@
6.3.2 Evaluation on Controlled Generation Scenes
@@
We conduct experiments on GPT-Neo-2.7B and GPT-J-6B, and find that $K_{IB}$ accurately predicts the classification accuracy. As depicted in Tab. 4, $K_{IB}$ correlates with classification accuracy with exactly the same trend. The interval-based method shows better IB metrics, which corresponds to its higher accuracy.
@@
Table 4: Explaining the superior performance of Inter in Inner-Probe using Mutual Information analysis on GPT-Neo-2.7B. Notably, the $K_{IB}$ results demonstrate a consistent trend with classification accuracy. Results of GPT-J-6B are not shown for space limit.
@@
6.4 Non-copyrighted Responses Filtering Module
@@
Incorporating intra-class copyright information improves performance, but using LSTM prior learnt from LLM has more to offer compared with using it during the unsupervised learning stage. As in Tab. 5, baseline [cho2022enhancing] with Setting1 improves accuracy by 3.94% and AUC by 0.108, yet still underperforms our method (Setting2) by an AUC of 0.104. Our method (Inter, Var and A-Var) shows consistent advantages and is more computationally efficient, as it doesn’t require fine-tuning on language models like BERT.
Table 5: Non-copyright Detection Accuracy and AUC Results.
Inter. Var. A-Var. Baseline [cho2022enhancing]
@@
6.5 Training Efficiency and Data Scale Effect
Robustness to Data Scale Large amounts of copyrighted data pose challenges to model training, necessitating a lightweight method. To evaluate data scaling effects, we test Inner-Probe by setting maximum token number $k$ to 7, and varying training samples from 160 to 1,600 and 12,800, with 6,400 inference samples. As shown in Tab. 6, training on 1,600 samples is sufficient for Inner-Probe to outperform baselines, while reducing training time (Tab. 7). The accuracy increases linearly with data scale and $k$. Our empirical study shows that while a small $k$ is generally sufficient, $k=7$ leads to model degradation.
Table 6: Data scale effect across different methods and models. $k$ denotes the sampling token number. The model accuracy strictly follows the increase in $k$ and data scale until $k=7$.
@@ 6.5 Training Efficiency and Data Scale Effect
Small (160,6400) 1 69.6 75.6 63.4 75.7 75.4 70.8 14.0 15.6 39.2 72.3
3 85.5 87.1 76.9 83.5 83.8 80.4
5 85.9 88.6 79.7 85.9 85.9 82.7
7 82.7 87.6 67.4 84.7 79.0 81.4
Medium (1600,6400) 1 88.6 89.0 89.9 91.2 89.5 88.3 14.0 15.6 42.7 79.4
3 92.8 93.1 91.5 92.6 91.8 91.1
5 93.3 93.6 92.3 93.2 93.1 91.3
7 93.4 94.0 92.5 93.2 93.0 91.7
Large (12800,6400) 1 92.0 93.3 93.4 93.9 92.2 93.3 14.0 15.6 42.8 79.9
3 94.8 95.4 94.3 95.3 94.3 94.4
5 95.4 95.5 94.3 95.4 94.9 94.7
7 95.2 95.7 94.8 95.3 95.2 95.1
Training Efficiency We conduct a comparative analysis of time cost across different methods and models. As shown in Tab. 7, Inner-Probe shows superior performance in less than 1.5 hours, far negligible than training or finetuning LLMs [Thoppilan2022LaMDALM].
Table 7: Training Efficiency Comparison
@@
6.6 Extended Discussion on Limitations
Our main experiments on Books3 and the Pile cover only a narrow slice of potential copyright data. Progress on dataset-level contribution analysis is further constrained by closed-source models, the scarcity of licensed copyright datasets, and unsettled legal laws. To mitigate these constraints, we complement the main experiments with a Matrix extension that expands coverage and improves transparency by pairing Matrix with the fully open-source LLM trained on it.
Extended experiments on Matrix Dataset
We evaluate six Matrix sub-datasets: Book-Education, Book-Finance, Book-History, Book-Math, Book-Medical, and Book-Law [zhang2024map]. These multilingual datasets better reflect broader and real-world attribution scenarios. Varying data scale, Inner-Probe attains at least 89% attribution accuracy and improves as the scale grows (Tab. 8). However, this means that for a specific dataset class containing extremely few samples, the performance degrades. In this case, a dedicated post-hoc detector can check outputs case by case [mitchell2023detectgpt, inan2023llama], since these rare samples may have limited influence on model behavior.
Effect of generation length
Contribution analysis accuracy slightly degrades when processing longer LLM outputs. In the continuation scene (as in Fig. 11(a)), overall accuracy is always around 75% but drops by about 3% on average as length increases from 256 to 512 (Tab. 9). Therefore, we recommend segmenting long LLM outputs into shorter ones before applying Inner-Probe. In all, Inner-Probe is accurate and scales with data on Matrix, but it weakens on very sparse classes and longer outputs; post-hoc checks and output segmentation are practical remedies.
Scope and outlook Even with the Matrix extension, important gaps remain. Closed-source models and non-public copyright datasets still limit end-to-end attribution, and standardized, licensed benchmarks are still emerging. Broader access to them will enable more comprehensive evaluations of contribution analysis methods.
Table 8: Extended Experiments on Map-Neo and Matrix Dataset. $k$ is set to 5 to show optimal performance.
Data scale $k=5$ Inter. Var. A-Var. Baseline1 Baseline2
Small (120,4800) 95.4 92.3 89.2 12.8 78.2
Medium (1200,4800) 97.9 97.5 97.6 12.8 86.7
Large (9600,4800) 99.8 99.2 99.0 12.8 89.5
Table 9: Contribution Analysis Accuracy with Varied Generation Length on Continuation Scene.
Generation Length Accuracy(%) 64 128 256 512
Inter. 73.65 72.95 77.82 71.73
A-Var. 72.95 72.96 80.56 78.94
Var. 70.98 77.03 75.15 72.98
7 Conclusion and Future Work
We highlight the need to identify specific copyrighted sub-datasets’ influence on LLM outputs. Our proposed framework Inner-Probe looks into LLM architectures, and discovers that MHA outputs provide more effective information than prompt-based, distance-based, or text classification methods. Information extraction, copyright sub-datasets contribution analysis and non-copyright responses filter modules are designed. Inner-Probe shows high accuracy and AUC comparable to text classification baselines when tested on original copyrighted texts. We further validate various generation scenarios, including prompt-induced book copyright infringements, continuation, copyright transfer, copyright elimination, and text mixture. Inner-Probe efficiently combines multi-modules for copyrighted data protection with minimal data, time, and GPU memory need, outperforming text classification, prompt-based and distance-based baselines by more than 15.02% accuracy and 0.104 AUC increase. Future work will extend to different fields such as Large Multimodal Models (LMMs) to verify its effectiveness in broader scenes.